OpenStudy (jkbo):

Please help Use the following scenario to answer the question: A bag contains 5 green marbles, 4 red marbles, 7 blue marbles, and 9 yellow marbles. If 3 marbles are drawn from the bag one at a time and not replaced, what is the probability of drawing a blue, yellow, and red marble in that particular order? A.63/3450 B.5/18 C.6/11 D.1/48

OpenStudy (anonymous):

B would be your best choice.

OpenStudy (jkbo):

@kelsey24 can you explain?

OpenStudy (john_es):

The probability of getting a blue marble is, \[P(blue)=\frac{7}{25}\] Now, you have to drop a ball from the 25. So in your bag now there are 24 balls. The probability of getting a yellow marble after the first blue extraction, \[P(yellow|\text{first blue})=\frac{9}{24}\] Now, your bag has 23 balls. The probability of getting a red marble if the first was blue and the second yellow, \[P(yellow|\text{first blue & second yellow})=\frac{4}{23}\] So the total probability of getting all of this at one time (equivalent to obtain it in three consecutive extractions without replacement), \[P=\frac{7}{25}\frac{9}{24}\frac{4}{23}=\frac{63}{3450}\]

OpenStudy (jkbo):

@john_ES I simplify it and I got 21/1150

OpenStudy (john_es):

Perfect.